# Your Turn: At Home Turning Angle Estimation for Parkinson’s Disease Severity Assessment

### v2

@@ 5 Experiments
Implementation and Evaluation – The experiment is conducted in PyTorch on one single NVIDIA 4060Ti GPU and a 12-core AMD Ryzen 5 5500 CPU. In our pipeline, the utilised FastPose model is trained on the MSCOCO pose estimation dataset [lin2014microsoft] and the Strided Transformer [stridedTransformer] is trained on Human3.6M [h36m_pami], following the standard set-up of related literature [martinez2017simple, zhang2022mixste]. These models are not optimised or fine-tuned to our free-living videos.
We evaluate our proposed method via three key metrics: accuracy, Mean Absolute Error (MAE) in degrees, and weighted precision (WPrec). Accuracy assesses the proportion of predicted angles that correctly fall into their respective bins, showing the categorical correctness of our predictions. MAE is calculated as the average of the absolute difference between the predicted values for angles, as well as speed in Turn-H3.6M, against ground truth. WPrec measures the percentage of true positive predictions for all positive predictions across angle bins, weighted by each bin’s sample size [rabby2023multi]. For example, if a turn is predicted as 90°, WPrec indicates the probability that the actual turn is 90°.
@@
Table 4: %$\%\text{Accuracy}_{\theta}$, $\text{MAE}_{\theta}$ () and %$\text{WPrec}_{\theta}$ of angle estimator against ground truth on Turn-REMAP. – There are a total of 1386 turns in Turn-REMAP. Accuracy is calculated by quantising the predicted angle into the nearest ${45}^{\circ}$ bin. $\text{MAE}_{\theta}$ shows the average error between the predicted continuous values and our labels. WPrec is the weighted average of each bin, with weights based on the sample size.
@@
[…]
Table 5: %$\%\text{Accuracy}_{\theta}$, $\text{MAE}_{\theta}$ (), $\text{MAE}_{\omega}$ (${}^{\circ}/\mathrm{s}$), and %$\text{WPrec}_{\theta}$ on Turn-H3.6M. – There are a total of 619 turns in Turn-H3.6M. Avg. represents the mean value across these groups. $\text{Avg}_{\omega}$ and $\text{Avg}_{d}$ represent the average turning speed and average duration within these groups.
@@ 6 Ablations
The accurate detection of 2D skeleton keypoints in each frame of our input clips is an important contributor to the overall accuracy of our method. Another fundamental concern is which single or combination of ‘body parts’ should be engaged for the computation of the turning angle. We investigate these two issues in our ablation study.
The effect of different 2D keypoints – We investigate how various 2D keypoint detectors impact the performance of the turning angle estimation in Turn-H3.6M. We applied SimplePose [xiao2018simple], HRNet [sun2019deep] and FastPose [fang2022alphapose] as prospective 2D keypoint detectors respectively and evaluated their performance in estimating turning angles. All three models were trained on the MS-COCO dataset [lin2014microsoft] following the same settings. HRNet and FastPose were chosen because they are state-of-the-art 2D keypoint detection models, while SimplePose, with its minimal yet effective design, was chosen to determine if more complex models are only overfitting the training dataset.
@@
On the human frontal plane, similar to knee and hip joints, shoulder joints are also potentially good indicators of the orientation of the body [lee1985determination]. However, PD patients have difficulty in maintaining lateral balance during weight shifts from one foot to the other while turning, demonstrating a greater inclination angle in the frontal plane than healthy controls [yang2016motion]. This suggests that, in PD, the shoulder joints may become less reliable for initiating turns, whereas combining the hip and knee joints shows less variability and may remain more stable in an upright stance. As shown in Table 7, the average predicted angle using both hip joints and knee joints yields the best accuracy, while averaging all three sets of joints gives the lowest MAE. for turning angle.
@@ 7 Discussion
Previous methods for turning analysis have been developed primarily for laboratory or clinical settings to evaluate scripted activities [shin2021quantitative, ribeiro2022public, lee2023gait, zeng2023video]. In Turn-REMAP, we record gait videos in a home-like, unobtrusive environment with PD and control subjects, and provide quantitative evaluations on the accuracy and estimation errors of turning angles during free-living activities. Pham et al. [Pham2017] also measured turning in a free-living environment, however, their method measured turning angles from IMUs alone, while our method is video-based. Pham et al. [Pham2017] recorded videos to manually validate their estimated results and report an overall error of $0.06^{\circ}$, but we contend that estimating turning angles at an accurate enough resolution to achieve such low error measurements by examining videos with the naked eye is unreliable. Some other IMU-based studies [El-Gohary2013, mancini2018turn, nouredanesh2021fall] have also extended their methodology to home environments, but none of these studies validated the measurement accuracy in the free-living setting.
@@
Although the overall measurement accuracy of our Turn-REMAP dataset is not yet robust enough for clinical diagnosis, it establishes a baseline for future, passive, video-based analysis of turning movements in indoor, free-living environments. Our manual examination of incorrectly classified video clips and their corresponding 3D skeletons revealed that depth reconstruction ambiguity is a significant factor [martinez2017simple, zhang2022mixste] that significantly affects the accurate calculation of turning angles. Recovering the missing depth from a 2D image is inherently an ill-posed problem as infinitely many 3D poses can project to the same 2D skeleton. Despite our utilised models being pretrained on large-scale laboratory 3D motion data, generalising this performance to reconstruct unseen poses in our in-the-wild PD turning dataset remains a substantial challenge.
@@
In Turn-REMAP, we find that the performance of our method on turns in free-living and loosely scripted activities is better than in clinical assessment (Table 4(a)). The reason for the performance degradation on these turns during the clinical assessment is the heightened occurrence of self-occlusion, where 40 of 49 of these turns are scripted 180°180°180  ° turns in a narrow hallway. This is confirmed by the findings in Table 4(b), which shows that locations like the dining room and hall, which have more occlusions, tend to have lower accuracy and higher MAE. Additionally, we find there is no significant difference in the performance of predicting turning angles for PD and Control (Table 4(c)), suggesting special PD gait characteristics do not affect the performance of our method. In contrast, IMU-based turning measurement methods [salarian2009analyzing, El-Gohary2013, mancini2018turn, haji2019turning] rely heavily on setting thresholds of angular velocity and relative orientation of the sensor attached to a single body part. Compared to skeleton-based models, the isolated sensory kinematic parameters are more easily affected by common PD symptoms, such as freezing of gait and slow turning speed [salarian2009analyzing].
(a) Comparison of distribution of predicted angle
(b) Comparison of distribution of MAE
Figure 5: Comparison of performance of the proposed model on Turn-REMAP and Turn-H3.6M. – In (a), mean predicted angles for Turn-REMAP are 76.1°76.1°76.1  °, 114.3°114.3°114.3  °, 141.4°141.4°141.4  °, and 152.6°152.6°152.6  ° at bins 90°90°90  °, 135°135°135  °, 180°180°180  °, and 225°225°225  °, respectively; for Turn-H3.6M, they are 84.4°84.4°84.4  °, 133.9°133.9°133.9  °, 178.2°178.2°178.2  °, and 207.0°207.0°207.0  °.
We further illustrate a comparison of the distribution of turns across different angle labels in the Turn-REMAP dataset against the distribution of turns at the same angles in the Turn-H3.6M dataset in Figure 5(a). The values in Turn-H3.6M are significantly closer to the expected bin angles, while in Turn-REMAP, the predicted angles tend to be underestimated. In these examined bins in Figure 5(a), the standard deviations for the Turn-REMAP dataset are 35.3°35.3°35.3  °, 37.6°37.6°37.6  °, 46.4°46.4°46.4  ° and 70.0°70.0°70.0  °, compared to those for the Turn-H3.6M dataset at 22.2°22.2°22.2  °, 20.5°20.5°20.5  °, 16.8°16.8°16.8  ° and 21.5°21.5°21.5  °, respectively. This shows that, compared to Turn-REMAP, there is less variability and uncertainty within each bin for predictions in Turn-H3.6M. The difference in performance is further shown in Figure 5(b), where we find that the distribution of MAE for the Turn-REMAP has a wider spread, while in Turn-H3.6M, 89.5% of the errors are smaller than 40°40°40  °.
@@
The statistics reveal challenges in generalising our pretrained human pose estimation model from the lab-based Turn-H3.6M dataset to the diverse, in-the-wild Turn-REMAP dataset. Different global position distributions [chai2023global], camera parameters [zhan2022ray3d], and diverse human body sizes and shapes, as well as articulated movements [gong2021poseaug, gholami2022adaptpose] highlight the need to enhance model robustness and adaptability to better handle real-world variability. To bridge this gap and enhance the ability to generalise to new, unseen data, it is crucial to implement domain adaptation strategies in deep learning models and conduct cross-dataset validation.
@@
Our model performs at 73.5% accuracy on Turn-H3.6M. This performance is limited by the inherent design of existing pose estimation algorithms, which are not specifically engineered to tackle biomechanical challenges, such as the analysis of turning characteristics. The training of these 2D-3D lifting models is usually guided by Mean Per Joint Position Error (MPJPE) [h36m_pami] loss, which focuses on minimising the absolute distance between the locations of the ground truth joints and the predictions. However, this criterion does not sufficiently address the requirements for temporal smoothness or accurate angular estimation. Therefore, further work on turning analysis involves building a downstream turning analysis algorithm based on the extracted deep learning features.
@@ 8 Conclusion and Future Work
While the accuracy of our models may not yet allow their application in the real world, they nevertheless establish a previously non-existent baseline and offer valuable insights for future video-based research in challenging free-living scenarios. Our sample size of 24 people, including 12 people with PD, demonstrates that our approach to detecting turning angles is promising and provides a proof of concept.
Automatically computing turning angles in a free-living environment is foundational for future longitudinal, in-home monitoring of PD. There are many potential avenues to build upon our work for more accurate turning angle estimation. Although Turn-REMAP and Turn-H3.6M only consist of trimmed turning clips, our methods can be extended to untrimmed videos. We could also infer additional turning metrics such as turning speed from the estimated turning angle. These metrics can be used to classify PD and control subjects, infer clinical rating scores of disease severity, or assess on/off medication status in free-living video recordings. Another extension for more accurate turning angle computation could be to replace our skeleton model with other models, such as via Human Mesh Recovery [goel2023humans] which could offer additional parameters for turning angle estimation.
@@ Acknowledgments
QC: Conceptualisation; Data curation; Formal analysis; Investigation; Methodology; Validation; Visualisation; Writing - original draft; and Writing - review & editing. CM: Resources; Data curation; writing - original draft preparation. Data curation; Formal analysis; Investigation; Supervision; Methodology; writing - original draft; and Writing - review & editing. AS: Conceptualisation; Data curation; Methodology. AM and AW: Supervision; Project administration. MM: Supervision; Project administration; Methodology; Writing - review & editing

### v3

@@ 5 Experiments
Implementation and Evaluation – Experiments are conducted in PyTorch on one single NVIDIA 4060Ti GPU and a 12-core AMD Ryzen 5 5500 CPU. We evaluate our proposed method via three key metrics: accuracy, Mean Absolute Error (MAE) in degrees, and weighted precision (WPrec). Accuracy assesses the proportion of predicted angles that correctly fall into their respective bins, showing the categorical correctness of our predictions. MAE is calculated as the average of the absolute difference between the predicted values for angles, as well as speed in Turn-H3.6M, against groundtruth. WPrec measures the percentage of true positive predictions for all positive predictions across angle bins, weighted by each bin’s sample size [rabby2023multi]. For example, if a turn is predicted as 90°, WPrec indicates the probability that the actual turn is 90°.
Pretraining Datasets– In our pipeline, the utilised FastPose model is trained on the MSCOCO [lin2014microsoft] pose estimation dataset, which provides multi-person keypoint annotations with 150K person instances in 57K images of complex everyday scenes containing common objects in their natural context. The Strided Transformer is trained on Human3.6M [h36m_pami], consisting of 3.6 million images featuring 7 healthy actors performing 15 everyday activities, as well as the 2D and 3D ground truth human joint locations using an accurate motion capture system in a controlled environment. The Strided Transformer is trained on subjects S1, S5, S6, S7 and S8 of Human3.6M, while subject S9 and S11 are held out for validation following the standard setup of related literature [martinez2017simple, zhang2022mixste]. Due to the constraints of an unobtrusive, free-living data collection environment, we did not perform any fine-tuning on Turn-REMAP. Instead, we investigated the generalization performance of the angular error across three data settings: 1) unseen test data from the challenging, real-world Turn-REMAP; 2) seen training data from subjects S1, S5, S6, S7 and S8 in Turn-H3.6M; and 3) unseen validation data from subjects S9 and S11 in Turn-H3.6M.
@@
Table 4: %$\%\text{Accuracy}_{\theta}$, $\text{MAE}_{\theta}$ () and %$\text{WPrec}_{\theta}$ of angle estimator against groundtruth on Turn-REMAP – There are a total of 1386 turns in Turn-REMAP. Accuracy is calculated by quantising the predicted angle into the nearest ${45}^{\circ}$ bin. $\text{MAE}_{\theta}$ shows the average error between the predicted continuous values and our labels. WPrec is the weighted average of each bin, with weights based on the sample size.
@@
Table 5: %$\%\text{Accuracy}_{\theta}$, $\text{MAE}_{\theta}$ (), $\text{MAE}_{\omega}$ (${}^{\circ}/\mathrm{s}$), and %$\text{WPrec}_{\theta}$ on Turn-H3.6M – There are a total of 619 turns in Turn-H3.6M. Avg. represents the mean value across these groups. $\text{Avg}_{\omega}$ and $\text{Avg}_{d}$ represent the average turning speed and average duration within these groups.
@@
[…]
Comparison between performance on Turn-REMAP and Turn-H3.6M – We further illustrate a comparison of the distribution of turns across different angle labels in the Turn-REMAP dataset against the distribution of turns at the same angles in the Turn-H3.6M dataset in Figure 5(a). The values in Turn-H3.6M are closer to the expected bin angles, while in Turn-REMAP, the predicted angles tend to be underestimated. In these examined bins labelled as 90°90°90  °, 135°135°135  °, 180°180°180  ° and 225°225°225  ° in Figure 5(a), the standard deviations for the Turn-REMAP dataset are 35.3°35.3°35.3  °, 37.6°37.6°37.6  °, 46.4°46.4°46.4  ° and 70.0°70.0°70.0  °, compared to those for the Turn-H3.6M dataset at 22.2°22.2°22.2  °, 20.5°20.5°20.5  °, 16.8°16.8°16.8  ° and 21.5°21.5°21.5  °, respectively. This shows that, compared to Turn-REMAP, there is less variability and uncertainty within each bin for predictions in Turn-H3.6M. The difference in performance is further shown in Figure 5(b), where we find that the distribution of MAE for the Turn-REMAP has a wider spread, while in Turn-H3.6M, 89.5% of the errors are smaller than 40°40°40  °.
(a) Comparison of distribution of predicted angle
(b) Comparison of distribution of MAE
Figure 5: Comparison of performance of the proposed model on Turn-REMAP and Turn-H3.6M – In (a), mean predicted angles for Turn-REMAP are 76.1°76.1°76.1  °, 114.3°114.3°114.3  °, 141.4°141.4°141.4  °, and 152.6°152.6°152.6  ° at bins 90°90°90  °, 135°135°135  °, 180°180°180  °, and 225°225°225  °, respectively; for Turn-H3.6M, they are 84.4°84.4°84.4  °, 133.9°133.9°133.9  °, 178.2°178.2°178.2  °, and 207.0°207.0°207.0  °.
Cross-sectional analysis of free-living turning characteristics – The turning angle serves as a fundamental metric for calculating more complex turning characteristics, such as maximum angular velocity. Here, we demonstrate the potential of the proposed method via a use case of PD classification from healthy controls. Since the turning angle is computed on a frame-by-frame basis, the maximum angular velocity $w_{max}$ of a turn can be calculated as
$w_{max}=\max(\theta_{0},\theta_{1},\ldots,\theta_{t},\ldots,\theta_{N-2})\times dt~{},$ (5)
where $dt$ represents the time interval for a single frame, calculated as the reciprocal of the frames per second (fps), and $\theta_{t}$ is the angle between two consecutive frames $t$ and $t+1$, where $t\in[0,N-2]$, and $N$ represents the number of frames, i.e.
$\theta_{t}=\frac{1}{2}\left(\sin^{-1}\left(\frac{||\mathcal{H}_{t}\times\mathcal{H}_{t+1}||}{||\mathcal{H}_{t}||\ ||\mathcal{H}_{t+1}||}\right)\ +\sin^{-1}\left(\frac{||\mathcal{K}_{t}\times\mathcal{K}_{t+1}||}{||\mathcal{K}_{t}||\ ||\mathcal{K}_{t+1}||}\right)\right)\ .$ (6)
[…]
Table 6: Group differences in turning angle and maximum angular velocity between PD and control subjects.
Measure Group Mean $\pm$ SD $t$-statistic ($p$) Effect Size (Cohen’s $d$)
Turning Angle () PD $92.65\pm 13.21$ $-1.73$ ($p=0.0998$) $-0.74$
Control $103.75\pm 16.75$
Max Angular Velocity (/s) PD $127.86\pm 29.77$ $-2.28$ ($p=0.0339$) $-0.97$
Control $160.19\pm 36.49$
@@ 6 Ablations
The accurate detection of 2D skeleton keypoints in each frame of our input clips is an significant contributor to the overall accuracy of our method. Another fundamental concern is which single or combination of ‘body parts’ should be engaged for the computation of the turning angle. We investigate these two issues in our ablation study.
The effect of different 2D keypoints – We investigate how various 2D keypoint detectors impact the performance of the turning angle estimation in Turn-H3.6M. We applied SimplePose [xiao2018simple], HRNet [sun2019deep] and FastPose [fang2022alphapose] as prospective 2D keypoint detectors respectively and evaluated their performance in estimating turning angles. All three models were trained on the MSCOCO dataset [lin2014microsoft] following the same settings. HRNet and FastPose were chosen because they are advanced 2D keypoint detection models, while SimplePose, with its minimal yet effective design, was chosen to determine if more complex models are only overfitting the training dataset.
@@
On the human frontal plane, similar to knee and hip joints, shoulder joints are also potentially good indicators of the orientation of the body [lee1985determination]. However, PD patients have difficulty in maintaining lateral balance during weight shifts from one foot to the other while turning, demonstrating a greater inclination angle in the frontal plane than healthy controls [yang2016motion]. This suggests that, in PD, the shoulder joints may become less reliable for initiating turns, whereas combining the hip and knee joints shows less variability and may remain more stable in an upright stance.
As shown in Table 8, the average predicted angle using both hip joints and knee joints yields the best accuracy, while averaging all three sets of joints gives the lowest MAE for turning angle.
@@ 7 Discussion
Previous methods for turning analysis have been developed primarily for laboratory or clinical settings to evaluate scripted activities [shin2021quantitative, ribeiro2022public, lee2023gait, zeng2023video]. In Turn-REMAP, we record gait videos in a home-like, unobtrusive environment with PD and control subjects, and provide quantitative evaluations on the accuracy and estimation errors of turning angles during free-living activities. Pham et al. [Pham2017] also measured turning in a free-living environment, however, their method measured turning angles from IMUs alone, while our method is video-based. Pham et al. [Pham2017] recorded videos to manually validate their estimated results and report an overall error of $0.06^{\circ}$, but we contend that estimating turning angles at an accurate enough resolution to achieve such low error measurements by examining videos with the naked eye is unreliable. Turning angle could also be calculated from 3D skeletons reconstructed using automatic methods [bogo2016keep, zhang2022mixste], whose performance is evaluated using Mean Per Joint Position Error (MPJPE) and Mean Per Joint Angle Error (MPJAE). Naturally, the granularity and resolution of the estimated turning angle depend on these performance metrics as well as the video’s frame rate. Though predicting the exact turning angle offers greater sensitivity, the increased precision often requires specialized setups or additional costs. In our free‑living scenario, automatically binning turning angles into discrete values can save clinicians significant time and effort in PD assessments — for example, consistently comparing 180° turns can help reveal early signs of disease progression in Parkinson’s disease [salarian2009analyzing]. Some other IMU-based studies [El-Gohary2013, mancini2018turn, nouredanesh2021fall] have also extended their methodology to home environments, but none of these studies validated the measurement accuracy in the free-living setting.
@@
Although the overall measurement accuracy of our Turn-REMAP dataset is not yet robust enough for clinical diagnosis, it establishes a baseline for future, passive, video-based analysis of turning movements in indoor, free-living environments. Our manual examination of incorrectly classified video clips and their corresponding 3D skeletons revealed that depth reconstruction ambiguity is an important factor [martinez2017simple, zhang2022mixste] that affects the accurate calculation of turning angles. Recovering the missing depth from a 2D image is inherently an ill-posed problem as infinitely many 3D poses can project to the same 2D skeleton. Despite our utilised models being pretrained on large-scale laboratory 3D motion data, generalising this performance to reconstruct unseen poses in our in-the-wild PD turning dataset remains a substantial challenge.
@@
In Turn-REMAP, we find that the performance of our method on turns in free-living and loosely scripted activities is better than in clinical assessment (Table 4(a)). The reason for the performance degradation on these turns during the clinical assessment is the heightened occurrence of self-occlusion, where 40 of 49 of these turns are scripted 180°180°180  ° turns in a narrow hallway. This is confirmed by the findings in Table 4(b), which shows that locations like the dining room and hall, which have more occlusions, tend to have lower accuracy and higher MAE. Additionally, we find there is no large difference in the performance of predicting turning angles for PD and Control (Table 4(c)), suggesting special PD gait characteristics do not affect the performance of our method. In contrast, IMU-based turning measurement methods [salarian2009analyzing, El-Gohary2013, mancini2018turn, haji2019turning] rely heavily on setting thresholds of angular velocity and relative orientation of the sensor attached to a single body part. Compared to skeleton-based models, the isolated sensory kinematic parameters are more easily affected by common PD symptoms, such as freezing of gait and slow turning speed [salarian2009analyzing]. However, a disadvantage of video-based methods is that, aside from occlusions—such as when the subject’s entire body is not in the line of sight—free-living scenarios may also introduce other challenges, including multi-person interactions. These could potentially be addressed using technology such as person re-identification for at-home health monitoring [zhu2015bridging, masullo2020person].
@@
The comparative statistics of the performance on Turn-REAMP and Turn-H3.6M reveal challenges in generalising our pretrained human pose estimation model from the lab-based Turn-H3.6M dataset to the diverse, in-the-wild Turn-REMAP dataset. Different global position distributions [chai2023global], camera parameters [zhan2022ray3d], and diverse human body sizes and shapes, as well as articulated movements [gong2021poseaug, gholami2022adaptpose] highlight the need to enhance model robustness and adaptability to better handle real-world variability. To bridge this gap and enhance the ability to generalise to new, unseen data, it is crucial to implement domain adaptation strategies in deep learning models and conduct cross-dataset validation. Although there is still room for improvement in the accuracy of turning angle estimation, our cross-sectional analysis reveals a statistically significant group difference between PD and C in maximum angular velocity. Specifically, PD subjects exhibited much slower turns and subtle reductions in turning angles, which is consistent with previous studies in different PD cohorts [El-Gohary2013, shah2020inertial]. This demonstrates that, even though our model was pretrained on public benchmark datasets composed of subjects without Parkinson’s disease, it was still able to characterize the impaired gait of older, early-stage PD subjects in Turn-REMAP.
@@
Our model performs at 73.5% (Table 7) accuracy on Turn-H3.6M. This performance is limited by the inherent design of existing pose estimation algorithms, which are not specifically engineered to tackle biomechanical challenges, such as the analysis of turning characteristics. The training of these 2D-3D lifting models is usually guided by Mean Per Joint Position Error (MPJPE) [h36m_pami] loss, which focuses on minimising the absolute distance between the locations of the groundtruth joints and the predictions. However, this criterion does not sufficiently address the requirements for temporal smoothness or accurate angular estimation. Therefore, further work on turning analysis involves building a downstream turning analysis algorithm based on the extracted deep learning features.
@@ 8 Conclusion and Future Work
While the accuracy of our models may not yet allow their application in the real world, they nevertheless establish a previously non-existent baseline and offer valuable insights for future video-based research in challenging free-living scenarios. Our sample size of 24 people, including 12 people with PD, demonstrates that our approach to detecting turning angles is promising and provides a proof of concept. Automatically computing turning angles in a free-living environment is foundational for future longitudinal, in-home monitoring of PD. There are many potential avenues to build upon our work for more accurate turning angle estimation. Although Turn-REMAP and Turn-H3.6M only consist of trimmed turning clips, our methods can be extended to untrimmed videos. We could also infer additional turning metrics such as turning speed from the estimated turning angle. These metrics can be used to classify PD and control subjects, infer clinical rating scores of disease severity, or assess on/off medication status in free-living video recordings. Another extension for more accurate turning angle computation could be to replace our skeleton model with other models, such as via Human Mesh Recovery [goel2023humans] which could offer additional parameters for turning angle estimation.
@@ Acknowledgments
QC: Conceptualisation; Data curation; Formal analysis; Investigation; Methodology; Validation; Visualisation; Writing - original draft; and Writing - review & editing. CM: Resources; Data curation; writing - original draft preparation. Data curation; Formal analysis; Investigation; Supervision; Methodology; writing - original draft; and Writing - review & editing. AS: Conceptualisation; Data curation; Methodology. AM and AW: Supervision; Project administration. MM: Supervision; Project administration; Methodology; Writing - review & editing.